ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

10. Partial Differential Equations

1991 MS Classification: 12, 31, 35, 39, 41, 65, 68, 76, 78, 80, 86, 92 Sastry, Seemanapalli V., Federal University of Paraiba, Brazil Application of Partial Differential Equations to Science and Engineering Partial differential equations in physical and geometrical problems were discussed. These may be time and space-dependent. Applications to science and engineering fields were presented. For Linear P. D. Equations in 2 independent variables, elliptic, parabolic and hyperbolic equations giving rise to Laplace, Heat Flow and Wave equations were suggested. Models of physical systems have been highlighted. Their utility in applied physics and engineering was demonstrated. Reference: Kreyszig, Erwin. Advanced Engineering Mathematics. John Wiley, New York. 1983, p. 988